Evaluate the ∫∫ y dS where S is the portion of the cylinder x2 + y2 = 3 that lies between z = 0 and z = 3.
1
Expert's answer
2020-04-22T18:38:13-0400

The cylinder is projected in a circle on the area Oxy. Use the polar coordinate system:

x = p cos ϕ

y = p sin ϕ

|J| = p

0 <= ϕ <= 2 π

Consider:

x2 + y2 = 3

p2 = 3

p = "\sqrt{3}"

0 <= p <= "\sqrt{3}"

"\iint ydS = \iint ydxdy = \int{_0^{2π}} dϕ \int{_0^{ \sqrt{3} }}p∙ p sin ϕ dp = 

\intop{_0^{2π}} sin ϕ∙" (p3/3)"\mid{_0^{ \sqrt{3}}}dϕ = 

\int{_0^{2π}} sin ϕ ∙ (3 \sqrt{3}/3) dϕ = \sqrt{3}∙ \intop{_0^{2π}}sin ϕ dϕ = \sqrt{3} ∙ (-cosϕ )

\mid{_0^{2π}} = - \sqrt{3}∙ (1-1) = 0"